# Triple Phase Transitions: Understanding the Learning Dynamics of Large Language Models from a Neuroscience Perspective

### v1

@@ A.4 Downstream Datasets
For MMLU, we use the original English dataset from Hendrycks et al. (2020) and its Japanese translation from OpenAI (2024). For control analysis, we use the Chinese translation from OpenAI (2024). Each of these datasets (English/Japanese/Chinese) comprises 13,571 samples. For CSQA, we use the original English dataset from Talmor et al. (2019) and the Japanese dataset from Kurihara et al. (2022), which contain 10,957 and 8,934 samples, respectively. For ARC (both the ARC-Challenge and ARC-Easy subsets), we use the original English dataset from Clark et al. (2018) and its Japanese translation, resulting in 7,778 samples for both the English and Japanese versions. For HellaSwag, we use the original English dataset from Zellers et al. (2019) and its Japanese translation, resulting in 9,658 samples for both the English and Japanese versions. We use the OpenAI API (GPT 4o-mini) for translation.

### v2

@@ A.4 Downstream datasets
MMLU assesses broad knowledge and problem-solving abilities using multidisciplinary coverage of 57 subjects, CSQA tests everyday conceptual commonsense reasoning, ARC probes elementary-level scientific knowledge, and HellaSwag assesses contextual commonsense reasoning in typical scenarios. For MMLU, we use the original English dataset from Hendrycks et al. (2020) and its Japanese translation from OpenAI (2024). For control analysis, we use the Chinese translation from OpenAI (2024). Each of these datasets (English/Japanese/Chinese) comprises 13,571 samples. For CSQA, we use the original English dataset from Talmor et al. (2019) and the Japanese dataset from Kurihara et al. (2022), which contain 10,957 and 8,934 samples, respectively. For ARC (both the ARC-Challenge and ARC-Easy subsets), we use the original English dataset from Clark et al. (2018) and its Japanese translation, resulting in 7,778 samples for both the English and Japanese versions. For HellaSwag, we use the original English dataset from Zellers et al. (2019) and its Japanese translation, resulting in 9,658 samples for both the English and Japanese versions. We use the OpenAI API (GPT 4o-mini) for translation.
In the probing analysis, each dataset is split into training and test datasets at a 4:1 ratio. Because MMLU comprises multiple subject areas, we split the dataset by subject. Furthermore, during the optimization of regularization parameters described in Section 3.3, to mitigate the subject-based bias of MMLU, we shuffle the training indices, and then perform cross-validation to ensure balanced distributions in each fold.